### v2

# Defining subsystems in Hilbert spaces with non-Euclidean metric

Abstract
This work outlines a consistent method of identifying subsystems in finite-dimensional Hilbert spaces, independent of the underlying inner-product structure. It has been well established that Hilbert spaces with modified inner-product, defined through the so-called metric operator, turn out to be the most natural ways to represent certain phenomena such as those involving balanced gain and loss resulting in pseudo-Hermitian Hamiltonians. For composite systems undergoing pseudo-Hermitian evolution, defining the subsystems is generally considered feasible only when the metric operator is chosen to have a tensor product form so that a partial trace operation can be well defined. In this work, we use arguments from algebraic quantum mechanics to show that the subsystems can be well-defined in every metric space – irrespective of whether or not the metric is of tensor product form. This is done by identifying subsystems with a decomposition of the underlying $C^{*}$-algebra into commuting sub-algebras. We show that different subsystem decompositions correspond to choosing different equivalence classes of the GNS representation. Furthermore, given a form of pseudo-Hermitian Hamiltonian, the choice of the Hamiltonian compatible metric characterizes the subsystem decomposition and as a consequence, the entanglement structure in the system. We clarify how each of the subsystems, defined this way, can be tomographically constructed and that these subsystems satisfy the no-signaling principle. With these results, we put all the choices of the metric operator on an equal footing.
I Introduction
It is an axiomatic assumption in quantum theory that the Hilbert space of a composite system has a tensor product decomposition. That is to say, a system with two degrees of freedom, represented by the set of variables $m_{1}$ and $m_{2}$ has the wavefunction $\psi(m_{1},m_{2})\in\mathcal{H}$ in the form $\psi(m_{1},m_{2})=\sum_{i}\psi_{i}(m_{1})\otimes\varphi_{i}(m_{2})\in\mathcal{H}_{1}\otimes\mathcal{H}_{2}$. Such a decomposition allows for a description of the sub-system localized to a sub-space of the total Hilbert space, with operations on this sub-system defined as operations on this sub-space. While this axiom has a footing in some fundamental principles of functional analysis and can be argued from deeper laws of nature  [2, 1], in certain situations, it does not hold good. The most celebrated demonstration of such a deviation is the system of indistinguishable particles [3, 4] and systems with infinite degrees of freedom described by quantum field theory [6, 5]. Another example, and the central topic of this work, is the system that evolves under pseudo-Hermitian Hamiltonians [7, 8]. Pseudo-Hermitian operators are the non-Hermitian operators with real eigenvalues. They arise naturally in certain optical settings with equal gain and loss [9, 10]. It was realized that such evolutions can be seen as unitary evolutions in Hilbert spaces with modified inner product structures induced by the metric operator  [12, 11]. In fact, such a description is necessary to avoid fallacious results like the violation of the no-signaling principle  [13]. For composite systems under such evolutions, the partial trace operation is well-defined if the metric operator allows a tensor product decomposition of the Hilbert space. Such choices of the metric operators are often chosen over other metric operators to define the subsystem  [14, 15]. However, it can so happen that for a given Hamiltonian, none of the choices of the metric operator allow for a tensor product structure [16]. In such situations, claiming that subsystems cannot be defined would be incorrect, especially if the experimentalist can probe the different degrees of freedom independently of each other. Therefore, to define subsystems in such Hilbert spaces one must use an operational definition of the subsystems.
In this work, we outline the problems that arise in defining a subsystem under pseudo-Hermitian evolution and resolve them using tools from the algebraic formulation of quantum mechanics. In the literature, the issue of defining subsystems independently of the Hilbert space structure has been explored by decomposing the underlying $C^{*}$-algebra [17, 18]. We use this algebraic basis to define subsystems in non-trivial inner product spaces. This allows us to put all the choices of the metric operator on equal footing. The central argument of this work is that the different decomposition of the underlying algebra amounts to choosing a particular family of inner product Hilbert spaces to represent a system. Consequently, while a pseudo-Hermitian evolution can be represented as unitary evolution in a family of inner product spaces, different choices of the inner product spaces can result in the probing of different sub-system decompositions. Finally, we show that the no-signaling principle is not broken under this description of the sub-system. While thus far the literature has relied on a genuine tensor product structure of the inner product spaces, we show that under any choice of the metric operator, the no-signaling principle holds for a well-defined sub-system.
I.1 Underlying vector space and the metric
The “pure” state of a classical system at a given time is given by a set of real numbers $\{m_{1},m_{2},...\}$ which correspond to a point in the configuration space. In other words, a measurement of the system can yield these values simultaneously. Let $m_{i}\in M_{i}$ where $M_{i}$ is a set of possible values the state can take. We call a system bipartite if the underlying configuration space is $M=M_{1}\times M_{2}$. This means that the experimentalist has the means to measure the parameters of $M_{1}$ and $M_{2}$ at the same time. For example, by placing the spin detectors at different points in space, one can measure a system’s spin and position. A quantum state is a complex-valued continuous function over the configuration space, given by $\psi(m_{1},m_{2},...)$. The space of all continuous functions on a locally compact Hausdaurff space $M$, denoted by $C^{\infty}(M)$, is a normed vector space that is complete with respect to the metric induced by the norm. This defines a Hilbert space. We are generally interested in a subspace $\mathcal{F}(M)\subset C^{\infty}(M)$, which is also complete with respect to the norm. The elements of this space are the function $f(m_{1},m_{2})$ such that $f(m_{1},\cdot)=f_{2}(\cdot)\in\mathcal{F}_{2}(M_{2})$ and $f(\cdot,m_{2})=f_{1}(\cdot)\in\mathcal{F}_{1}(M_{1})$, where $\mathcal{F}_{1}(M_{1})$ and $\mathcal{F}_{2}(M_{2})$ are set of continuous functions over $M_{1}$ and $M_{2}$. Let the set $\mathcal{F}_{1}(M_{1})\otimes\mathcal{F}_{2}(M_{2})$ denote the closure of the set of linear combinations of functions of the form $\varphi(m_{1},m_{2})=f_{1}(m_{1})\times f_{2}(m_{2})$. Then, we have the equality $\mathcal{F}_{1,2}(M_{1}\times M_{2})=\mathcal{F}_{1}(M_{1})\otimes\mathcal{F}_{2}(M_{2})$ if at least one of the sets $\mathcal{F}_{i}(M_{i})$ is a complete normed vector space   [19].
Therefore, for a system whose underlying configuration space has a Cartesian product form, we assume that the vector space of pure quantum states has a tensor product decomposition. This vector space is converted to a Hilbert space by introducing an inner product map. As we discuss below, the tensor product decomposition of the resulting Hilbert space depends on one’s choice of the inner product map.
I.2 Overview of quantum mechanics in non-Euclidean metric space
Given two elements $\psi$ and $\phi$ on a vector space $\mathcal{V}$, define a Euclidean (or trivial) inner product map $\langle\psi|\phi\rangle\ =\bar{\psi}\phi$, where $\bar{\psi}$ is the conjugate transpose of the element $\psi$. With respect to this inner product, one can define a class of non-Euclidean (or non-trivial) inner products facilitated by a Hermitian, invertible operator $G$ called a metric operator: $\langle\psi|\phi\rangle_{G}\ =\bar{\psi}G\phi=\langle\psi|G\phi\rangle$. We therefore have a class of inner product spaces $\mathcal{H}_{G}:=(\mathcal{V},G)$, where we will denote $\mathcal{H}_{\mathbb{1}}$ simply as $\mathcal{H}$. It can be shown that if $\mathcal{H}_{\mathbb{1}}$ is a Hilbert space, then $\mathcal{H}_{G}$ is also a Hilbert space, i.e. complete with respect to the norm induced by their respective inner product map. The issues in defining subsystems in these Hilbert spaces arise since for an underlying vector space $\mathcal{V}_{1}\otimes\mathcal{V}_{2}$, one has the freedom to endow it with a metric operator that is not in the form of $G_{1}\otimes G_{2}$. In such a case, the partial trace operation on $\mathcal{H}_{G}$ is not well-defined.
An observable is defined as an operator with real expectation values. In $\mathcal{H}_{G}$, such operators must satisfy the relation $G\mathcal{O}=\mathcal{O}^{\dagger}G$, called pseudo-Hermiticity  [8]. An operator can be an observable in two different Hilbert spaces since the metric satisfying the pseudo-Hermiticity condition is not unique. Let $\mathcal{H}_{G^{\prime}}$ be such a Hilbert space. It is characterized by an isometric map $T:\mathcal{H}_{G^{\prime}}\to\mathcal{H}_{G}$ which satisfies the relation $\eta^{\prime}T^{-1}=U\eta$ where $U$ is any unitary map ($U^{\dagger}U=1$). Therefore the isometries $T$ are parameterized by the operator $U$, hence we will denote them by $T_{U}$. The metric operators are related by $T_{U}^{\dagger}GT_{U}=G^{\prime}$, where this relation is independent of $U$. The operator $\mathcal{O}$ is an observable in both metric spaces if and only if $[T_{U},\mathcal{O}]=0$. Given metric operators $G$ and $G^{\prime}$, this relation holds for a unique unitary $U$.
A pseudo-Hermitian operator in $\mathcal{H}_{G}$ can be mapped to a Hermitian operator in $\mathcal{H}$ via $\mathcal{O}\rightarrow\mathcal{O}_{\eta}=\eta\mathcal{O}\eta^{-1}$, where $\eta$ is the unique positive square root of the metric operator $G$. Note that $\eta:\psi\rightarrow\eta\psi$ is an isometry between the spaces $\mathcal{H}_{G}$ and $\mathcal{H}$. Furthermore, since $G^{\prime}$ is another possible metric operator, $\mathcal{O}_{\eta^{\prime}}=\eta^{\prime}\mathcal{O}\eta^{\prime-1}$ is another Hermitisization of the observable $\mathcal{O}$. These two Hermitian forms of $\mathcal{O}$ are unitarily equivalent: $\mathcal{O}_{\eta}=U\mathcal{O}_{\eta^{\prime}}U^{\dagger}$. In fact, one can show that $[T_{U},\mathcal{O}]=0$ if and only if $\mathcal{O}_{\eta}=U\mathcal{O}_{\eta^{\prime}}U^{\dagger}$[16] (see Fig.  1). It can be shown that the space of all allowed metric operators compatible with Hamiltonian $H$, can be parametrized by the set of real vectors $\vec{\lambda}$[11]. Given $\{\psi_{i}\}$ and $\{\phi_{j}\}$ the left and right eigenvectors of $H$ respectively, (given it has a non-degenerate spectrum), the most general metric is given by $G_{\lambda}=\sum_{i}\lambda_{i}|\psi_{i}\rangle\langle\phi_{i}|$.
Figure 1: Isometric maps between different metric Hilbert spaces. The state $\psi$ in $\mathcal{H}_{G^{\prime}}$ can be mapped to a state in $\mathcal{H}$ using isometric maps $T_{U}$. The resulting states are related by a unitary $U$ in $\mathcal{H}$.
I.3 Open-system perspective
Pseudo-Hermitian evolutions in quantum systems have been experimentally exhibited many times in recent years  [20, 21, 22]. This is typically achieved by introducing an interacting potential wherein the amplitude and frequency of the gain and loss of the probability current are carefully controlled to maintain the potential’s PT-symmetry  [23]. That is, the interacting Hamiltonian is invariant under the action of the parity and time reversal operations. Within certain parameters (referred to as the ”unbroken” PT-symmetry regime) the Hamiltonian has a Pseudo-Hermitian form  [24]. Due to the presence of interaction terms in the Hamiltonian, such an evolution is sometimes seen as open system dynamics, even though the effective Hamiltonian has real eigenvalues. A quintessential feature that separates a pseudo-Hermitian evolution from a genuinely non-Hermitian (non-unitary) evolution is the revival of complete information of the system throughout the dynamics, measured as probability values or the information back-flow. It is argued  [25, 26] that such a complete revival is due to the presence of a finite-sized entangled partner in the environment of the system, where the Hermitisization operation $\psi\rightarrow\eta\psi$ is seen as an embedding of the system into a larger Hermitian system. This description, while illuminating, views pseudo-Hermiticity as a non-trivial extension of quantum mechanics. As we will elucidate in the subsequent sections, the pseudo-Hermitian Hamiltonian, although simulated through interaction terms, induces a unitary evolution.
A large class of open system dynamics follows a Lindblad-type master equation, which can be recast as an evolution under an effective non-Hermitian Hamiltonian and a “jump-term”. Post-selecting on the trajectories without a jump, the evolution of the system can be cast as an evolution under the non-Hermitian Hamiltonian. If the effective Hamiltonian is pseudo-Hermitian, then one can show that the Lindblad equation is simply the von Neumann equation for a redefined pseudo-Hermitian state, and the jump term vanishes (see A for details), effectively reproducing the unitary dynamics.
II The C*-algebra
The equivalence between the trivial and non-trivial metric space representation of quantum mechanics becomes clear if we look at it from an algebraic point of view. In this section, we will see how the same quantum system can have both trivial and non-trivial metric space descriptions (representations), and that these descriptions are unitarily equivalent. This viewpoint helps us in the subsequent subsection to define a consistent definition of a subsystem that is independent of the tensor product decomposition of the Hilbert space.
We start with a C*-algebra of operators $\mathcal{A}$ and a positive linear functional called a state $\omega$ (with norm 1) defined on this algebra. The algebra and the state $(\mathcal{A},\omega)$ together represent the statistical data of the given quantum experiment, e.g. $A\in\mathcal{A}$ can represent the measurements that can be performed on a system, and the number $\omega(A)$ will then represent the result of such operation. A faithful representation of the algebra is an *-isomorphism to the set of bounded operators on a Hilbert space $\pi:\mathcal{A}\rightarrow\pi(\mathcal{A})\subset\mathcal{B}(\mathcal{H})$. This is possible since the set of bounded operators on a complex Hilbert space is also a C*-algebra and the star operation is naturally defined via $\pi(A)^{*}=\pi(A^{*})$. The GNS construction  [27], is the construction of the Hilbert space description of this system, which involves first constructing a $C^{*}$ algebra on which the state is faithful, i.e., has a trivial kernel. This algebra $\mathcal{U}_{\omega}$ is constructed by first identifying the kernel of the state $K_{\omega}=\{X\in\mathcal{A}\,\text{ s.t. }\,\omega(X^{*}X)=0\}$, and “quotienting out” the kernel from the algebra: $\mathcal{U}_{\omega}:=\mathcal{A}/K_{\omega}$. In other words, the elements of $\mathcal{U}_{\omega}$ consist of the equivalence classes $\psi_{A}=\{A+X\,\text{ s.t. }\,X\in K_{\omega}\}$. One can see that the functional $\omega$ is faithful on $\mathcal{U}_{\omega}$. Moreover, $\mathcal{U}_{\omega}$ is a complex vector space, a property one can check by using the fact that $AX\in K_{\omega}\,\forall\,X\in K_{\omega}$, i.e. $K_{\omega}$ is a left ideal of $\mathcal{A}$. In the next step of the GNS construction, the state functional $\omega$ is used to define an inner product on the vector space $\mathcal{U}_{\omega}$ via $\langle\psi_{A}|\psi_{B}\rangle=\omega(A^{*}B)$. With this added structure, $\mathcal{U}_{\omega}$ becomes a pre-Hilbert space which can be completed to form a Hilbert space $\mathcal{H}_{\omega}=\overline{\mathcal{U}}_{\omega}$. The bar over the algebra denotes the completion with respect to the norm induced by the inner product. Therefore, the set of vectors $\{\psi_{A}\}$ is dense in $\mathcal{H}_{\omega}$. If one defines a representation given by the following action on the states: $\pi_{\omega}(A)\psi_{B}=\psi_{AB}$, one can then derive the relation $\omega(A^{*}B)=\langle\psi_{\mathbb{1}}|\pi_{\omega}(A^{*}B)\psi_{\mathbb{1}}\rangle$, where $\mathbb{1}$ is the identity element of the algebra $\mathcal{A}$. Therefore, $\psi_{\mathbb{1}}$ is a cyclic vector of the representation $(\mathcal{H}_{\omega},\pi_{\omega})$. This construction gives us a faithful cyclic representation of the algebra $\mathcal{U}_{\omega}$, with the representation space $\mathcal{H}_{\omega}$, the map $\pi_{\omega}$, and the cyclic vector $\psi_{\mathbb{1}}$, such that $\omega(A)=\langle\psi_{\mathbb{1}}|\pi(A)\psi_{\mathbb{1}}\rangle_{\mathcal{H_{\omega}}}$.
It is also a well-known fact that such a representation is unique up to a unitary equivalence. In the subsequent part of this article, we will assume that given a $C^{*}$ algebra, one can always construct a faithful cyclic representation of a system $(\mathcal{A}_{\omega},\omega)$. The state $\omega$ is associated with an operator $\rho_{\omega}$ such that $\omega(A)=tr(\rho_{\omega}\pi_{\omega}(A))$. In particular, we will demand that the operator $\rho_{\omega}$ is an element of the set $\pi_{\omega}(\mathcal{A})$, in which case it is a unique operator fixed by the faithful representation map $\pi_{\omega}$.
II.1 Cyclic representations in different metric spaces
Consider two faithful cyclic representations of the system $(\mathcal{A},\omega)$: (¯$(\bar{\mathcal{H}}_{\omega},\bar{\pi}_{\omega},\bar{\Phi}_{\omega})$ and $(\mathcal{H}_{\omega},\pi_{\omega},\Phi_{\omega})$. Let us assume that the two Hilbert spaces are different metric spaces where $\mathcal{H}_{\omega}$ is endowed with metric $G$ and ¯$\bar{\mathcal{H}}_{\omega}$ endowed with metric $G^{\prime}$. Therefore $\mathcal{H}_{\omega}=\mathcal{H}_{G}$ and ¯$\bar{\mathcal{H}}_{\omega}=\mathcal{H}_{G^{\prime}}$. Since these are cyclic representations, they must be unitarily equivalent, hence there exists a unitary map [^1] By “unitary” we mean an invertible isometry defined everywhere on the Hilbert space. In the most general sense dual of a unitary is equal to its inverse $T_{U}:\mathcal{H}_{G^{\prime}}\rightarrow\mathcal{H}_{G}$ between the two Hilbert spaces whose action is given by $\Psi_{A}\rightarrow T_{U}\Psi_{A}$, where $T_{U}$ is a bounded and positive operator. Therefore we should have $\langle\psi_{A}|\psi_{A}\rangle_{\mathcal{H}_{G^{\prime}}}=\langle T_{U}\psi_{A}|T_{U}\psi_{A}\rangle_{\mathcal{H}_{G}}=\langle\psi_{A}|T_{U}^{*}T_{U}\psi_{A}\rangle_{\mathcal{H}_{G^{\prime}}}$, or in other words $T_{U}^{*}T_{U}=1$. One can check that the dual operator $T_{U}^{*}$ is given by the relation $T_{U}^{*}=G^{\prime-1}T^{\dagger}G$. Combining this with the unitarity property, we have the relation $G^{\prime}=T_{U}^{\dagger}GT_{U}$. We had discussed this relation in the introduction of the metric formalism. Here we have seen that this is a consequence of the the unitary equivalence of the cyclic representation of the same underlying algebra. The operator $T_{U}$ intertwines the algebras $\pi_{G}(\mathcal{A})$ and $\pi_{G^{\prime}}(\mathcal{A})$ such that $\pi_{G}(A)=T_{U}\pi_{G^{\prime}}(A)T_{U}^{-1}\,\forall\,A\in\,\mathcal{A}$. Note that the Hermitisization operation $\eta:\mathcal{H}_{G}\rightarrow\mathcal{H}:\psi\rightarrow\eta\psi$ is a special case of such a unitary operation. Therefore, given a system $(\mathcal{A},\omega)$, it can have unitarily equivalent faithful cyclic representations in different metric Hilbert spaces and always has a trivial metric (Hermitian) description  (see also  [28]).
To summarize, by defining the unitary operation as an invertible-isometric map between two Hilbert spaces, possibly of different metric structures, it becomes clear that the unique equivalence class of the cyclic representation of the state contains its not-trivial metric space representations.
II.2 Partitions of a system
Let us have a system $(\mathcal{A},\omega)$ with cyclic representation $(H_{\omega},\pi_{\omega},\Omega_{\omega})$ and two sub-algebras $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$. Consider the cyclic representations of the two subsystems $\{\mathcal{A}_{1},\omega_{1}\}$ and $\{\mathcal{A}_{2},\omega_{2}\}$ given by $\{\mathcal{H}_{\omega_{1}},\pi^{1}_{\omega_{1}},\Omega_{\omega_{1}}\}$ and $\{\mathcal{H}_{\omega_{2}},\pi^{2}_{\omega_{2}},\Omega_{\omega_{2}}\}$ respectively, where $\omega_{1}$ and $\omega_{2}$ are the restriction of $\omega$ on $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ respectively. We define the Kernel space of these states as $K_{\omega_{1}}=\{X\in\mathcal{A}_{1}\,\text{and}\,\omega(X^{*}X)=0\}$ and similarly $K_{\omega_{2}}$, and use them to construct the Hilbert spaces $\mathcal{H}_{\omega_{1}}=\overline{\mathcal{A}_{1}/K_{\omega_{1}}}$ and $\mathcal{H}_{\omega_{2}}=\overline{\mathcal{A}_{2}/K_{\omega_{2}}}$. Define the product of these spaces $\mathcal{H}_{\omega_{1}}\odot\mathcal{H}_{\omega_{2}}$ as a linear combination of the elements of the form $\psi_{A_{i}}^{1}\psi_{B_{j}}^{2}$[29]. Here we of course have $A_{i}\in\mathcal{A}_{1}$ and $B_{j}\in\mathcal{A}_{2}$. This forms a linear vector space with an inner product defined through the state $\omega(A_{i}B_{j})=\omega_{1}(A_{i})\omega_{2}(B_{j})$. Note that this is not a faithful state over $\mathcal{H}_{\omega_{1}}\odot\mathcal{H}_{\omega_{2}}$. This is because the kernel of $\omega$ cannot be constructed solely from the kernels $K_{\omega_{1}}$ and $K_{\omega_{2}}$. One therefore defines a Hilbert space called the direct product Hilbert space by quotienting out this kernel: $\mathcal{H}_{\omega_{1}}\otimes\mathcal{H}_{\omega_{2}}=(\overline{\mathcal{H}_{\omega_{1}}\odot\mathcal{H}_{\omega_{2}})/K_{\omega}}$. Under this equivalence, the elements of the form $\psi_{A_{i}}^{1}\psi_{B_{j}}^{2}$ are mapped to the elements of the form $\psi_{A_{i}}^{1}\otimes\psi_{B_{j}}^{2}\in\mathcal{H}_{\omega_{1}}\otimes\mathcal{H}_{\omega_{2}}$. Finally, $(\mathcal{H}_{\omega_{1}}\otimes\mathcal{H}_{\omega_{2}},\pi^{1}_{\omega_{1}}\otimes\pi^{2}_{\omega_{2}},\Omega_{\omega_{1}}\otimes\Omega_{\omega_{2}})$ is a cyclic representation of the algebra $\mathcal{A}_{1}\vee\mathcal{A}_{2}$ defined as the minimal algebra containing $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$. Note that if $\pi_{\omega}(A_{i}B_{j})=\pi^{1}_{\omega_{1}}(A_{i})\otimes\pi^{2}_{\omega_{2}}(B_{j})$, one can map the elements of $\mathcal{H}_{\omega_{1}}\otimes\mathcal{H}_{\omega_{2}}$ to the elements of $\mathcal{H}_{\omega}$. However, if all the elements of the larger algebra $\mathcal{A}$ can be written as a linear combination of terms like $A_{i}B_{j}$, this mapping becomes an isomorphism between $\mathcal{H}_{\omega_{1}}\otimes\mathcal{H}_{\omega_{2}}$ and $\mathcal{H}_{\omega}$. This is possible only if the elements of the two sub-algebras commute with each other and the linear combination of the product of their elements is equal to the total algebra. This condition, purely at the level of algebras, is a necessary condition for the factorization of the Hilbert space $\mathcal{H}_{\omega}\cong\mathcal{H}_{\omega_{1}}\otimes\mathcal{H}_{\omega_{2}}$, which we regard as the conventional partitioning of a quantum system. However, the algebraic statement provides a more operational meaning to the partitioning of a system and is useful in situations where Hilbert space does not factorize for a multipartite system. Therefore, motivated by earlier works  [17], we provide the following definition for a bi-partition of a quantum system (generalization to a multi-partition is straightforward).
Definition 1.
Bi-partition of a system: Given a system $(\mathcal{A},\omega)$, a bi-partition of this system is the pair $\{(\mathcal{A}_{1},\omega),(\mathcal{A}_{2},\omega)\}$ if the following conditions are satisfied
  1. 1.
    The sub-algebras $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ commute: $\forall\,A_{i}\in\mathcal{A}_{1}$ and $B_{j}\in\mathcal{A}_{2}$, we have $[A_{i},B_{j}]=0$.
  2. 2.
    The minimal algebra generated by their union is the algebra $\mathcal{A}$, a relation denoted by the equation $\mathcal{A}_{1}\vee\mathcal{A}_{2}=\mathcal{A}$. $\forall\,X\in\mathcal{A}$, $X=\sum_{i,j}A_{i}B_{j}$ where $A_{i}\in\mathcal{A}_{1}$ and $B_{j}\in\mathcal{A}_{2}$. Note that this form of a general element of $\mathcal{A}_{1}\vee\mathcal{A}_{2}$ is due to the commutativity of these algebras.
Note that the $C^{*}$-subalgebras $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ can be replaced by the commuting observable algebras $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$. These are the so-called Jordan-Lie-Banach algebras, whose complexification gives us the corresponding $C^{*}$-algebra  [30].
Continuing the above discussion, let us suppose that a particular cyclic representation of $(\mathcal{A},\omega)$, given by $(\mathcal{H}_{\omega},\pi_{\omega},\Omega_{\omega})$ allows a genuine bi-partition of the system into $(\mathcal{A}_{1},\omega)$ and $(\mathcal{A}_{2},\omega)$ and has the metric operator $G$, such that $\mathcal{H}_{\omega}=\mathcal{H}_{G}=\mathcal{H}_{\omega}^{1}\otimes\mathcal{H}_{\omega}^{2}$. One therefore concludes that $G$ has a tensor product structure $G=G_{1}\otimes G_{2}$ where $G_{1}$ and $G_{2}$ are the metric operators for Hilbert spaces $\mathcal{H}_{\omega}^{1}$ and $\mathcal{H}_{\omega}^{2}$, respectively. The corresponding Hermitisization map from such a Hilbert space will maintain the tensor product structure: $\eta=\eta_{1}\otimes\eta_{2}:\mathcal{H}_{G}\rightarrow\mathcal{H}_{\mathbb{1}}=\mathcal{H}^{1}\otimes\mathcal{H}^{2}$. Therefore, a genuine bi-partition in any metric space is equivalent to a genuine bi-partition in the trivial metric Hilbert space. Hence, we will restrict ourselves to genuine bi-partitions in the trivial metric Hilbert space.
Let us now restrict ourselves to defining subsystems in representations $(\mathcal{H}_{G},\pi_{G})$, where $G\neq G_{1}\otimes G_{2}$, while we do have $\mathcal{H}_{\mathbb{1}}=\mathcal{H}^{1}\otimes\mathcal{H}^{2}$, i.e. the bi-partition in the non-trivial metric Hilbert space is not a genuine bi-partition. An observable $\pi_{G}(A)$ under this representation, is also an observable in the Hilbert space $\mathcal{H}_{G^{\prime}}$ if $\pi_{G}(A)$ commutes with the unitary map $T_{U}:\mathcal{H}_{G^{\prime}}\rightarrow\mathcal{H}_{G}$. Therefore, given an observable, there can be an infinite number of metric spaces to choose from. However, if we fix a certain set of observables $\{\pi_{G}(A_{i})\}_{i}$ to probe the system, such that the only operator that commutes with all the elements of the set is a multiple of identity, then the metric space under which this set remains a set of observables is fixed  [31]. Such a set is called an irreducible set of operators as they do not have any common invariant subspace. This discussion points to the relationship between the choice of the metric operator and the choice of observables. The question then naturally arises, given a partitioning of an algebra, to what extent is the metric operator fixed? The following theorem answers this question.
Theorem 1.
For a system $\{\mathcal{A},\omega\}$, any two cyclic representations $\pi$ and $\pi^{\prime}$, with representation spaces $\mathcal{H}_{G}$ and $\mathcal{H}_{G^{\prime}}$ respectively, correspond to the same bi-partitioning into system 1 and 2 if and only if they belong to the equivalence class given by:
$[\pi]\equiv\{\pi^{\prime}(\cdot)=T_{U}^{-1}\pi(\cdot)T_{U};\,\eta^{\prime}T_{U}^{-1}=(U_{1}\otimes U_{2})\eta\}.$ (1)
Here $U_{1}$ and $U_{2}$ are unitary operators corresponding to the two subsystems.
Proof.
Let $\mathcal{A}_{\pi}^{1}=\{A\in\mathcal{A}:\eta\pi(A)\eta^{-1}=\mathcal{O}_{A}\otimes\mathbb{1}\}$ and $\mathcal{A}_{\pi^{\prime}}^{1}=\{A\in\mathcal{A}:\eta^{\prime}\pi_{G^{\prime}}(A)\eta^{\prime-1}=\mathcal{O}_{A}^{\prime}\otimes\mathbb{1}\}$ be the two sub-algebras defined w.r.t. two different metric operators $G=\eta^{2}$ and $G^{\prime}={\eta^{\prime}}^{2}$ respectively. One defines the complementary sub-algebras $\mathcal{A}_{\pi}^{2}=\{A\in\mathcal{A}:\eta\pi_{G}(A)\eta^{-1}=\mathbb{1}\otimes\mathcal{O}_{A}\}$ and similarly, $\mathcal{A}_{\pi^{\prime}}^{2}$. This gives us two bi-partitions of the algebra $\mathcal{A}$ into $\mathcal{A}^{1}_{\pi}\vee\mathcal{A}^{2}_{\pi}$ and $\mathcal{A}^{1}_{\pi^{\prime}}\vee\mathcal{A}^{2}_{\pi^{\prime}}$.
Now we will determine the condition for the equality of the two algebras $\mathcal{A}_{\pi}^{1}\sim\mathcal{A}_{\pi^{\prime}}^{1}$. Let $X\in\mathcal{A}_{\pi}^{1}$. We know that the operators under the two Hilbert spaces are related by $\pi^{\prime}(X)=T_{U}^{-1}\pi(X)T_{U}$ such that $\eta^{\prime}T_{U}^{-1}=U\eta$. Combining these two relations we have $\eta^{\prime}\pi_{G^{\prime}}(X)\eta^{\prime-1}=U\eta\pi_{G}(X)\eta^{-1}U^{\dagger}$. Therefore $X\in\mathcal{A}_{1}^{\prime}$ if and only if $U\mathcal{O}_{X}\otimes\mathbb{1}U^{\dagger}=\mathcal{O}_{X}^{\prime}\otimes\mathbb{1}$ for the given unitary $U$. Similarly we have $\mathcal{A}_{\pi}^{2}\sim\mathcal{A}_{\pi^{\prime}}^{2}$ if and only if $U(\mathbb{1}\otimes\mathcal{O}_{Y})U^{\dagger}=\mathbb{1}\otimes\mathcal{O}_{Y}^{\prime}$ for all $Y\in\mathcal{A}^{2}_{\pi^{\prime}}$. Unitaries satisfying these conditions must always be of the form $U=U_{1}\otimes U_{2}$ (see e.g. [32, 33]). Therefore, the two partitions induced by the representations $\pi$ and $\pi^{\prime}$ are equivalent if and only if the interwining isometry $T_{U}$ satisfies the relation $\eta^{\prime}T_{U}=U_{1}\otimes U_{2}\eta$. It is straightforward to see that representations related by such an intertwiner form an equivalence class. ∎
In a typical experiment, the pseudo-Hermitian Hamiltonian is simulated through post-selection or equal gain-loss mechanisms. The form of the Hamiltonian is fixed by the experimentalist, and therefore one is restricted to a class of metric spaces compatible with this Hamiltonian. Such metric spaces are fixed by the isometry $T_{U}$, or equivalently, the unitary $U$. As a simple corollary of the above theorem, the partitioning of the system is then fixed by the choice of the metric space.
Corollary 1.1.
Given a pseudo-Hermitian operator $H$ such that $H^{\dagger}G=GH$, the choices of the compatible representation spaces $\mathcal{H}_{G^{\prime}}$ with equivalent partitioning are those belonging to the equivalence class [G] defined by
$[G]\equiv\{G^{\prime}=T_{U}^{\dagger}GT_{U};\,\eta^{\prime}T_{U}^{-1}=(U_{1}\otimes U_{2})\eta\,\text{ and }\ [T_{U},H]=0\}.$ (2)
Proof.
Given two cyclic representations $(\mathcal{H}_{G},\pi)$ and $(\mathcal{H}_{G^{\prime}},\pi^{\prime})$, such that $H=\pi(X)=\pi^{\prime}(X)$, therefore the representation spaces are $H$-compatible. There exists an isometry $T_{U}$ between the Hilbert spaces such that $\pi=T_{U}\pi^{\prime}T_{U}^{-1}$ which implies $[\pi(X),T_{U}]=0$. From the previous theorem, the two sub-algebra decompositions $\mathcal{A}_{\pi}^{1}\vee\mathcal{A}_{\pi}^{2}$ and $\mathcal{A}_{\pi^{\prime}}^{1}\vee\mathcal{A}_{\pi^{\prime}}^{2}$ are equivalent if and only if $\eta^{\prime}T_{U}^{-1}=(U_{1}\otimes U_{2})\eta$. ∎
As a consequence of the above results, we see that while the choice of the metric does not influence the system’s properties, viz. the expectation values of the observables, it can change the properties of the subsystem, e.g. the entanglement between the subsystems. This is not unlike in the case of global unitary rotation on the system, which is simply a choice of a global reference frame, which does not change the expectation values on the system, but does change the subsystems’ properties like the entanglement between them. On the other hand, a local unitary transformation keeps the subsystem properties intact. The above theorem explains this observation from an algebraic point of view: performing a local unitary transformation in the trivial metric Hilbert space is the same as choosing a metric space from an equivalence class of metric operators, and therefore keeps the underlying algebraic partitioning the same.
Tensor product structure on $\mathcal{H}_{G}$: A tensor product structure (TPS)  [17] on a Hilbert space $\mathcal{H}_{G}$ is an isomorphism $\varphi:\mathcal{H}_{G}\rightarrow\mathcal{H}=\mathcal{H}_{1}\otimes\mathcal{H}_{2}$ where $\dim(\mathcal{H}_{1})\times\dim(\mathcal{H}_{2})=\dim(\mathcal{H}_{G})$. We know that this isomorphism given by the map  $\varphi:\,|\psi\rangle=\eta|\psi\rangle$. Consider another Hilbert space $\mathcal{H}_{G^{\prime}}$ with the metric $G^{\prime}$ and a TPS $\varphi^{\prime}:\mathcal{H}_{G^{\prime}}\rightarrow\mathcal{H}=\mathcal{H}_{1}\otimes\mathcal{H}_{2}:\psi\rightarrow\eta^{\prime}\psi$. We have also seen that there is an isometry between the two Hilbert spaces, as defined in previous sections, is $\mathcal{T}:\mathcal{H}_{G^{\prime}}\rightarrow\mathcal{H}_{G}:\,\psi\rightarrow T_{U}|\psi\rangle$. Therefore, the map $\varphi^{\prime}\circ\mathcal{T}^{-1}:\psi\rightarrow\eta^{\prime}T_{U}^{-1}\psi$ defines another TPS over the Hilbert space $\mathcal{H}_{G}$. This gives us a family of TPS on $\mathcal{H}_{G}$ parametrized by the operator $\eta^{\prime}$, or equivalently by $U$. If the unitaries parameterizing the TPS are local transformations, i.e. of the form $U=U_{1}\otimes U_{2}$, we say they generate equivalent TPS [^2] Following the terminology of  [17]. From corollary 1.1 we see that:
Remark: a family of equivalent TPS on $\mathcal{H}_{G}$ corresponds to the equivalence class of metric operators $[G]\equiv\{G^{\prime}=T_{U}^{\dagger}GT_{U};\,\eta^{\prime}T_{U}^{-1}=(U_{A}\otimes U_{B})\eta\},$ and consequently the same partitioning of the total algebra.
II.3 Partial trace
Given a genuine bi-partition $\mathcal{H}=\mathcal{H}_{1}\otimes\mathcal{H}_{2}$, we have the isomorphism $\pi(\mathcal{A}_{1})\cong\mathcal{B}(\mathcal{H}_{1})$. Therefore, the subsystem can be represented as a bounded operator on $\mathcal{H}_{1}$ in the following way: $\omega(A)|_{A\in\mathcal{A}_{1}}=tr(\rho_{\omega}\mathcal{O}_{A}\otimes\mathbb{1}_{2})=tr(\rho_{\omega}^{r}\mathcal{O}_{1}).$ The partial trace map $\rho_{\omega}\rightarrow\rho_{\omega}^{r}$ is a completely positive and trace-preserving map. From an operational point of view, the subsystem is completely defined by the algebra $\mathcal{A}_{1}$ and the value of the functional $\omega$ on this algebra. Therefore, in representations where such a genuine bi-partition is not possible, although the local description of the subsystem and operations do not exist, the subsystem is well defined. In this section, we will discuss how to define the subsystems in a non-trivial metric space and propose an extended notion of partial trace.
Earlier we had discussed the irreducibility of a set of operators. A representation of a quantum system $(\mathcal{A},\omega)$ given by the map $\pi_{\omega}$ is called an irreducible representation if the set of operators $\pi(\mathcal{A})$ is an irreducible set. Irreducibility of a representation, it turns out, depends entirely on the purity of the state $\omega$. A state is called a pure state if it can not be written as a convex combination of other states.
With these discussions, it becomes clear that the reduced state of a pure state can be a mixed state (not pure): Given a pure state $(\mathcal{A},\omega)$, the restriction of the state to a sub-algebra $\mathcal{A}_{1}$ can satisfy the condition $\omega(A)=\sum\lambda_{\alpha}\omega^{\alpha}(A),\ \forall A\in\mathcal{A}_{1}$, where $\omega^{\alpha}$ are pure states. Of course, we know that for genuine bi-partitions the partial trace of a pure state can often result in a mixed state. For a state $\omega=\sum\lambda_{\alpha}\omega_{\alpha}$, the density matrix representation under the $\pi_{\omega}$ is $\rho_{\omega}=\lambda_{\alpha}\rho_{\alpha}$, where $\rho_{\alpha}$ are pure states.
The entropy of the sub-system An important result in C-* algebra then says that a cyclic representation of $(\mathcal{A},\omega)$ is irreducible if and only if the state $\omega$ is a pure state over $\mathcal{A}$. The representation of a mixed state $(\mathcal{A},\omega)$ is therefore reducible into invariant sub-spaces $\mathcal{H}_{\omega}^{k}$ : $\pi_{\omega}=\oplus_{k}\ \pi_{\omega}^{k}$, and the Hilbert space decomposes into the invariant sub-spaces $\mathcal{H}_{\omega}=\oplus_{k}\ \mathcal{H}_{\omega}^{k}$. This decomposition can be used to construct a spectral decomposition of the state, where the invariant sub-spaces are the eigen-sub-spaces of the state: $\rho_{\omega}=\oplus e_{k}\rho_{k}$. Here $e_{k}$ are the eigenvalues of the density matrix, and if the eigen-spectrum has no degeneracy, then $\rho_{k}$ are pure states. This corresponds to the preparation of the state $\omega=\sum e_{k}\omega_{k}$. The entropy of this preparation, which we define as the Shanon entropy of the vector $\vec{e}$: $H(\vec{e})=-\sum{e_{k}\ln e_{k}}$ is therefore equal to the von-Newmann entropy of the corresponding density matrix $S_{VN}=\rho_{\omega}\ln(\rho_{\omega})$. Complications that arise with the degeneracies in the spectrum have been discussed in the appendix B (see also [34, 35, felski]).
Consider a special case, as discussed in  [37], where the subsystem of interest is the invariant subspace of the representation. Given a Hilbert space of the form $\mathcal{H}=\oplus_{\alpha}\ \mathcal{H}_{\alpha}$ with the metric $G=\oplus_{\alpha}\ G_{\alpha}$, consider a state $\rho_{\omega}=\oplus_{\alpha}\lambda_{\alpha}{\rho_{\omega}}^{\alpha}$. If all $G_{\alpha}$ operators were the same, a partial trace can be given by the density metric $\rho_{\omega}^{r}=\sum_{\alpha}\lambda_{\alpha}\rho_{\omega}^{\alpha}\,\in\mathcal{H}_{\alpha}$. However, such a summation of density matrices is not well defined when they belong to different metric spaces. In fact, one is required to perform a parallel transport of the operators to another metric space in order to add them. The isometry $T_{U}$ defined earlier provides such a metric-compatible parallel transport of operators. For this example, one can define the reduced state $\rho^{r}(\beta)=\sum_{\alpha}\lambda_{\alpha}T_{\alpha,\beta}\ {\rho_{\omega}}^{\alpha}{T_{\alpha,\beta}}^{-1}$ on the Hilbert space $\mathcal{H}_{\beta}$. Here, $T_{\alpha,\beta}$ is an isometric operation between the Hilbert spaces $\mathcal{H}_{\alpha}\to\mathcal{H}_{\beta}$ and can be written as $T_{\alpha,\beta}=\eta_{\beta}^{-1}U\eta_{\alpha}$ for a unitary operator $U$. This procedure defines a unique CPTP map between $\mathcal{B}(\oplus\mathcal{H}_{\alpha})\rightarrow\mathcal{B}(\mathcal{H}_{\beta})$ and is a consistent definition of a partial trace. Note that the reduced state thus defined is independent of the unitary operators $U$ or equivalently the choice of isometries $T_{\alpha,\beta}$. Indeed, if one chooses $U=\mathbb{1}$ for all $T_{\alpha,\beta}$, the partial trace is compatible with the partial trace by Hermitisization: $\eta_{\beta}\rho^{r}(\beta)\eta^{-1}_{\beta}=tr_{X}(\eta\rho_{\omega}\eta^{-1})$.
III Tomography and the no-signaling principle
A tomographically complete set of observables, given by $\{A_{i}\}\subset\mathcal{A}$, satisfies the condition: $\omega_{1}(A_{i})=\omega_{2}(A_{i})\iff\omega_{1}=\omega_{2}$ for all states $\omega_{1}$ and $\omega_{2}$ on $\mathcal{A}$. Therefore, such a set completely characterizes any given state $\omega$. We will show that the corresponding density matrices, $\bar{\rho}\in\mathcal{H}_{G}$, and $\rho_{\eta}\in\mathcal{H}$ are tomographically equivalent i.e., based on the distribution of $\omega$ on a tomographically complete set of observables, the experimentalist can construct either of the states. The density matrix for the state, under the representation $(\mathcal{H},\pi)$, is constructed through a sampling operator $W$[38] which is given as
[…]
Here $\mu$ is a probability measure on the set of observables $\pi(\mathcal{A})=\mathcal{B}(\mathcal{H})$ and $\langle\mathcal{O}_{i}|\,\rho\rangle=tr(O_{i}^{\dagger}\rho)=\omega(A_{i})$ is the Hilbert-Schmidt inner product between the operators $\mathcal{O}_{i}=\pi(A_{i})$ and $\rho$. Faithfully reconstructing the state requires that the operator $W$ be full rank, which is a constraint on the measure $\mu$. It is said to be a tomographically complete measure if $W$ is full rank. Note that given a tomographically complete set of observables $\{\mathcal{O}_{i}\}$ in the Hilbert space $\mathcal{H}$, the set of observables $\{\bar{\mathcal{O}}_{i}\}$, where $\bar{\mathcal{O}_{i}}=\eta^{-1}\mathcal{O}_{i}\eta$, is a tomographically complete set in $\mathcal{H}_{G}$. Furthermore, since the norm in the two Hilbert spaces is related by $||\mathcal{O}||=||\bar{\mathcal{O}}||_{G}$, the measure in the space $\mathcal{B}(\mathcal{H}_{G})$, defined by $d\mu_{G}(\bar{\mathcal{O}})=d\mu(\mathcal{O})$, is also a probability measure. Combining these observations we can construct the sampling operator in $\mathcal{B}(\mathcal{H}_{G})$: $\overline{W}=\eta^{-1}W\eta$. Therefore, it becomes clear that if the measure $\mu$ gives a tomographically complete set of observables in $\mathcal{B}(\mathcal{H})$, then $\mu_{G}$ is tomographically complete. Furthermore, if $\mu$ is a tight frame (i.e. $W=\mathbb{1}$), then $\mu_{G}$ is also a tight frame.
In an experiment, the knowledge of a state is in terms of the statistical data of the observable (e.g. $A_{i}$) and the probability of the value of the measurement being $m$, denoted by $\{m,p(m|A_{i})\}$. The connection between this data and the equation (3) is given by Born’s rule which tells us that the inner product $\langle\pi(A_{i})|\rho)\rangle$ is the expectation value of the measurement results of the observable $A_{i}$: $\sum_{m}mp(m|A_{i})=\langle\pi(A_{i})|\rho\rangle=\omega(A_{i})$. From our discussion in the previous sections, we know that this quantity is independent of the chosen cyclic representation of the system $(\mathcal{A},\omega)$ i.e. $\langle\mathcal{O}|\,\rho\rangle=\langle\bar{\mathcal{O}}|\,\bar{\rho}\rangle$.
Take the example of the spin 1/2 system. The set $\{\mathbb{1},\vec{\sigma}\}$ forms a tight frame in the trivial metric Hilbert space $\mathcal{H}=\mathbb{C}^{2}$. This is also the minimal group that induces an irreducible representation on $\mathcal{H}$[39]. The tight frame consists of measurement of the spin in three orthogonal directions since the Pauli matrices form the orthogonal basis for the space of observables in two dimensions: $tr(\sigma_{i}\sigma_{j})=2\delta_{ij}$. A general observable is therefore given by $\vec{\sigma}.\hat{n}$ and is associated with a direction $\hat{n}$ in 3-dimensional space. Let $p(m|\hat{n}_{i})$ be the probability of getting the outcome $m$, given the spin was measured in the $\hat{n}_{i}$ direction, corresponding to the measurement of the $\sigma_{i}$ observable. To make this measurement we use a Stern-Gerlach set-up where the particles of a specific velocity pass through the magnetic field $\vec{B}=B\hat{n}_{i}^{\perp}$ (see the appendix C for details) and then through a magnetic field in $z$-direction which splits the beam in two parts corresponding to the values of $m\in\{\pm\dfrac{1}{2}\}$. By controlling the direction and magnitude of the magnetic field $\vec{B}$, one produces the experimental data and constructs the density matrix: $\rho=\mathbb{1}/2+\sum_{m,i}mp(m|\hat{n}_{i})\sigma_{i}$. Note that while associating the operator $\sigma_{i}$ with the direction $\hat{n}_{i}$, the experimentalist makes two choices. The first is the choice of a tight frame from the family of unitarily equivalent tight frames. Since the set of operators $\{\mathbb{1},U\sigma_{i}U^{\dagger}\}$ also forms a tight frame in $\mathcal{H}$, the density matrix thus constructed is fixed only up to a unitary equivalence. The second choice is the metric space in which the operators $\{\sigma_{i}\}$ and their unitary equivalents are observables. One can equivalently choose the set of operators $\{\bar{\sigma}_{i}=\eta^{-1}\sigma_{i}\eta\}$ which are the observables in the metric space $\mathcal{H}_{G}$ with metric $G=\eta^{2}$. They also form a tight frame with $tr(\bar{\sigma}^{*}_{i}\bar{\sigma}_{j})=2\delta_{ij}$. With this choice of the metric space, the resultant matrix $\bar{\rho}=\mathbb{1}/2+\sum_{m,i}mp(m|n_{i})\bar{\sigma}_{i}$ is a density matrix in $\mathcal{H}_{G}$.
Now consider a composite system of two distinguishable spin-1/2 particles. In $\mathcal{H}=\mathcal{H}_{1}\otimes\mathcal{H}_{2}$ the set of operators $\{\mathbb{1},\sigma_{i}\otimes\sigma_{j},\mathbb{1}\otimes\sigma_{i},\sigma_{i}\otimes\mathbb{1}\}$ forms the tight frame. The experimental setup consists of two independent Stern-Gerlach apparatus mentioned above with independent control over the magnetic field $\vec{B}_{1}$ and $\vec{B}_{2}$. We then collect the data to construct the probabilities $p(m_{1},m_{2}|\hat{n}^{1}_{i},\hat{n}^{2}_{j})$ of obtaining the result $m_{1}$ as a result of measurement in $\hat{n}^{1}_{i}$ direction on particle 1 and $m_{2}$ as a result of measurement in $\hat{n}^{2}_{j}$ direction on particle 2. Born’s rule gives us the trace $\sum m_{1}m_{2}p(m_{1},m_{2}|\hat{n}^{1}_{i},\hat{n}^{2}_{j})=tr(\rho_{12}\sigma_{i}\otimes\sigma_{j})$. If the two particles are accessible to two distinct parties, namely Alice and Bob, then Alice only has the data $\{m_{1},p(m_{1}|\hat{n}^{1}_{i})\}$ which is a subset of the complete set of data of the system. When can we say that this subset of the complete data describes a subsystem? From the point of view of Alice, there has been no difference between the previous experiment and this experiment, and therefore she constructs a state in $\mathcal{H}_{1}$ namely $\rho_{1}$ using the Born rule $\sum m_{1}p(m_{1}|\hat{n}^{1}_{i})=tr(\rho_{1}\sigma_{i})$. From the global point of view, this expectation value is represented as $\sum m_{1}p(m_{1}|\hat{n}^{1}_{i})=tr(\rho_{12}\sigma_{i}\otimes\mathbb{1})$. ($\mathbb{1}$ corresponds to switching off the secondary magnetic field in the S-G set-up). Therefore, Alice’s statistical data corresponds to the set of observables $\{\sigma_{i}\otimes\mathbb{1}\}$, which forms a sub-algebra. With $\{\mathbb{1}\otimes\sigma_{i}\}$ forming the complementary sub-algebra, one can say that this data completely characterizes the sub-system, in this case, represented by the density matrix $\rho_{1}$.
Consider the non-trivial metric representation of the system of two particles in $\mathcal{H}_{G}$ as the state $\bar{\rho}_{12}$. Alice, still in possession of the data $\{m,p(m,\hat{n}^{1}_{i})\}$, might be aware of the global metric structure $G$, but due to the non-separability of the metric, she does not have a notion of local observables. Therefore, she cannot construct a density matrix for particle 1 alone which can be seen as a reduced state of $\bar{\rho}_{12}$. Nevertheless, from the global point of view, we have $\sum m_{1}p(m_{1}|\hat{n}^{1}_{i})=tr(\bar{\rho}_{12}\eta^{-1}(\sigma_{i}\otimes\mathbb{1})\eta)$. This is because we have $tr(\bar{\sigma}_{j}^{*}\bar{\sigma}_{j})=2\delta_{ij}$, which means that these measurements correspond to measurements in orthogonal directions. Therefore, Alice’s data also corresponds to the set of observables $\{\eta^{-1}(\sigma_{i}\otimes\mathbb{1})\eta\}$, which forms a sub-algebra. With $\{\eta^{-1}(\mathbb{1}\otimes\sigma_{i})\eta\}$ forming the complementary sub-algebra together generating the total $C^{*}$ algebra $\mathcal{B}(\mathcal{H}_{G})$, Alice’s data corresponds to a well-defined subsystem of $\mathcal{H}_{G}$. But as we have established, the data $\{m,p(m|\hat{n}_{i}^{1})\}$ is faithfully represented by $\rho_{1}\in\mathcal{H}_{1}$, therefore, it captures a well-defined subsystem of $\mathcal{H}_{G}$. These arguments can be generalized for an arbitrary finite-dimensional system, and therefore, we have the following theorem:
Theorem 2.
Given a density matrix $\bar{\rho}_{12}$ in $\mathcal{H}_{G}$, the density matrix $\rho_{1}=tr_{2}(\eta\bar{\rho}_{12}\eta^{-1}$) represents a well-defined sub-system of the system represented by $\bar{\rho}_{12}$.
The no-signaling principle: The no-signaling principle states that the statistics of the subsystem measured by Alice denoted by $\{m_{1},p(m_{1}|\hat{n}^{1}_{i})\}$ should not change by an operation on another subsystem belonging to Bob. One can show that this principle holds in all metric spaces. If Alice’s and Bob’s subsystems correspond to the sub-algebra $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ respectively, then for their shared system $\bar{\rho}_{12}$ in the metric space $(\mathcal{H}_{G},\pi_{G})$, such that $\eta\pi_{G}(B)\eta^{-1}=\mathbb{1}\otimes\mathcal{O}_{B}$, we have the following relation:
$\rho_{1}=tr_{2}(\eta\bar{\rho}_{12}\eta^{-1})=tr_{2}(\eta\pi_{G}(B)\bar{\rho}_{12}\eta^{-1}),\,\forall\,B\in\mathcal{A}_{2}.$ (4)
The second equality comes from the no-signaling principle in trivial metric space: $tr_{2}(\eta\pi_{G}(B)\bar{\rho}_{12}\eta^{-1})=tr_{2}(\mathbb{1}\otimes\mathcal{O}_{B}\rho_{12})=tr_{2}(\rho_{12})$. Since we have already established that the state $\rho_{1}$ faithfully represents Alice’s subsystem, we can conclude that the no-signaling principle for a bi-partition, as given in the definition 1, is obeyed in every metric space. In the literature [15], one typically shows that the no-signaling principle is satisfied when the metric is separable $G=G_{1}\otimes G_{2}$, which is only a specific case of the more general principle that we have exhibited here through equation (4).
IV Conclusion
The axiom of the tensor product decomposition of the Hilbert space of composite quantum systems restricts how we understand a subsystem. In a vast family of quantum systems, such a decomposition is not possible, and consequently, a partial trace operation is not well defined. This issue arises in the non-trivial metric space representation of composite systems, which forms a natural description of the systems undergoing balanced gain and loss under the pseudo-Hermitian Hamiltonian. In this work, we have used algebraic tools to provide an operational definition of subsystems in such Hilbert spaces with non-trivial metric structures. We show that given an algebra and a state over it, one can equivalently construct a Hermitian or a pseudo-Hermitian cyclic representation, which are unitarily equivalent if we use the term unitary in its most general sense. While given a pseudo-Hermitian Hamiltonian, one can choose from a family of Hamiltonian compatible metric spaces, we show that different choices of the metric operator can correspond to choosing different system decompositions. Since choosing a metric is equivalent to choosing a reference frame for the system, it makes sense that choosing two different metric representations changes the subs-system properties. Furthermore, the choice of sub-system decomposition is equivalent to a choice of an equivalence class of metric spaces. Finally, we show that from an experimentalist’s point of view, local tomography on a sub-system undergoing pseudo-Hermitian evolution creates no paradoxes, and allows one to construct a local description of the subsystem in trivial metric space. The no-signaling principle, which so far has been restricted to tensor product decomposable Hilbert spaces, can be consistently shown to hold good for an algebraically defined sub-system, even when the Hilbert spaces do not have a genuine tensor product structure.
After the mathematical description of pseudo-Hermitian quantum mechanics as a trivial extension of normal quantum mechanics was established, most studies dealing with composite systems in this field have been wary of using the non-separable metric structure due to the ill-defined partial trace operation in such scenarios. From this work, it is clear that such metric operators are as physically meaningful as the separable ones, and there is no clear physical reasoning for choosing a separable metric operator over the non-separable operators to define the subsystems. We hope that this work will be helpful in a better understanding of composite quantum systems evolving under the $\mathcal{P}\mathcal{T}$-symmetric non-Hermitian Hamiltonians.
Acknowledgments: H.B. thanks A P Balachandran for useful discussions and Namit Anand for his valuable comments on the draft.
Appendix A Lindbladian dynamics in $\mathcal{H}_{G}$
Given the Lindblad dynamics of the state $\rho$
$\dot{\rho}(t)=-i[H,\rho(t)]+\sum_{j}-\dfrac{1}{2}\{L_{j}^{\dagger}L_{j},\rho(t)\}+L_{j}\rho L_{j}^{\dagger},$ (5)
define $H_{e}=H-i\sum_{i}L_{j}^{\dagger}L_{j}$ (effective non-Hermitian Hamiltonian) so that the above equation is can be recast in the following form
$\dot{\rho}(t)=-i(H_{e}\rho(t)-\rho(t)H_{e}^{\dagger})+\sum_{j}L_{j}\rho(t)L_{j}^{\dagger}.$ (6)
This shows that the Lindblad equation can be written as an evolution under a non-Hermitian Hamiltonian and an extra term called a “jump term”. Define $M_{j}=L_{j}\sqrt{dt}$ and $M_{0}=\mathbb{1}-iH_{e}dt$, where $dt$ is an infinitesimal amount of time. Then the following infinitesimal CP map corresponds to the above Lindblad equation:
$\Lambda_{t,t+dt}(\rho(t))=\rho(t+dt)=M_{0}\rho(t)M_{0}^{\dagger}+\sum_{j}M_{j}\rho(t)M_{j}^{\dagger}.$ (7)
The probability of the jump in this infinitesimal time interval $t$ to $t+dt$ is given by $dp=\sum_{j}tr(L_{j}^{\dagger}L\rho(t))dt$ and the probability of no jump in this interval is $p_{0}=1-dp=tr(M_{0}^{\dagger}M_{0}\rho(t))$.
If the effective Hamiltonian is pseudo-Hermitian ($H_{e}^{\dagger}\eta^{2}=\eta^{2}H_{e}$), then, the transformation $\rho\rightarrow\rho\eta^{2}=\bar{\rho}$ rewrites the Lindblad equation as
$\dot{\bar{\rho}}(t)=-i[H_{e},\bar{\rho}(t)]+\sum_{j}L_{j}\bar{\rho}(t)L_{j}^{*}$ (8)
Here $A^{*}:=\eta^{-2}A^{\dagger}\eta^{2}$. This corresponds to the following stochastic evolution in $\mathcal{H}_{G}$
$\bar{\rho}(t+dt)=\bar{\Lambda}_{t,t+dt}(\bar{\rho}(t))=M_{0}\bar{\rho}(t)M_{0}^{*}+\sum_{j}M_{j}\bar{\rho}(t)M_{j}^{*}.$ (9)
Now the jump probability within the time interval $t$ to $t+dt$ is $dp=\sum_{j}tr(M_{j}^{*}M_{j}\bar{\rho})=\sum_{j}tr(L_{j}^{\dagger}\eta^{2}L_{j}\rho)$ and the probability of no jump is $p_{0}=tr(M_{0}^{*}M_{0}\bar{\rho})$. However, one can check that $M_{0}^{*}M_{0}=\mathbb{1}$, therefore, $p_{0}=tr(\bar{\rho})$. Since one can always choose $\eta^{2}$ s.t. $tr(\bar{\rho})=1$, the pseudo-Hermiticity condition tells us $\sum_{j}tr(L_{j}^{\dagger}L_{j}\eta^{2})=0$. From here we can conclude $\sum_{j}tr(L_{j}^{\dagger}\eta^{2}L_{j}\rho)=0$ In other words, the jump probability in the redefined Lindbladian equation vanishes if the effective Hamiltonian is Hermitian, indicating a unitary evolution.
Appendix B Ambiguities in entropy of the state
Given a finite-dimensional representation of the C*-algebra $\mathcal{A}$, given by $(\mathcal{H},\pi)$, a state $\omega$ on this algebra is uniquely represented by the element $\rho_{\omega}\in\pi(\mathcal{A})$ of this algebra, such that
$\omega(A)=tr(\rho_{\omega}\pi(A))\hskip 28.45274pt\forall\ A\in\mathcal{A}.$
Now, if the state $\omega$ is pure, so is the density matrix $\rho_{\omega}$. If it is a mixed state, it can be written as a convex combination of other pure states: $\omega=\sum_{\alpha}\lambda_{\alpha}\omega^{\alpha}$, and the corresponding density matrix can be written as $\rho_{\omega}=\sum_{\alpha}\lambda_{\alpha}{\rho_{\omega}}^{\alpha}$ where $\rho_{\omega}^{\alpha}$. are pure state density matrices Furthermore, the representation is reducible into a direct sum of irreducible representations $\rho_{\omega}=\oplus_{i}p_{i}\rho_{i}$. Note that here $\rho_{i}$ need not be pure states. The convex sum decomposition of the state $\omega$ is not unique since the state space is not a simplex. Therefore, a state can be prepared from an ensemble of pure states in many ways, represented by the probability vector $\vec{\lambda}$ s.t. $\sum_{\alpha}\lambda_{\alpha}=1$. The entropy associated with such a preparation is given by the Shannon entropy of the probability vector $\vec{\lambda}$: $H(\vec{\lambda})=-\sum_{\alpha}\lambda_{\alpha}\ln(\lambda_{\alpha})$ The von-Newmann entropy of the density matrix $\rho_{\omega}$, on the other hand, is given by $S(\rho_{\omega})=-tr(\rho_{\omega}\ln\rho_{\omega})=\sum_{i}e_{i}\ln e_{i}$ where $\vec{e}$ is the spectrum of the state $\rho_{\omega}$. A density matrix can be written as a convex combination of pure density matrices in many non-unique ways: $\rho=\sum_{i}\lambda_{i}|\psi_{i}\rangle\langle\psi_{i}|$. It turns out that the Shannon entropy of all such preparations (given by $H(\vec{\lambda})$ ) is bounded below by the von-Newmann entropy of the state: $S(\rho)=\inf\{H(\vec{\lambda});\rho=\lambda_{i}|\psi_{i}\rangle\langle\psi_{i}|\}$. Similarly one can construct the unique entropy of the state $\omega$ on the algebra $\mathcal{A}$ as the infimum of all the preparations of $\omega$: $S(\omega)=\inf\{H(\vec{\lambda});\omega=\sum_{\alpha}\lambda_{\alpha}\omega^{\alpha}\}$. We therefore have a unique notion of the entropy of a state if $S(\omega)=S(\rho_{\omega})$. As shown in [34], this is true only under faithful representations with no multiplicities. The most general relation between the two entropies is given as follows:
$S(\rho_{\omega})=S(\omega)+\sum_{i}p_{i}\ln m_{i}$ (10)
where $m_{i}$ are the multiplicities of the $i-th$ irreducible subspace and $p_{i}$ are the weights of each invariant subspace in the representation $\rho_{\omega}=\oplus_{i}p_{i}\rho_{i}$. In this work, we will define the entropy of the state $\omega$ as the von Neumann entropy of the unique density matrix $\rho_{\omega}$, implying that we will restrict to the multiplicity-free faithful representation.
Appendix C Spin tomography
General spin tomography: Given eigenvectors $|m\rangle$ of $\sigma_{z}$ with eigenvalues $m\in\{\dfrac{1}{2},-\dfrac{1}{2}\}$, the probability of obtaining the result $m$ after measuring the spin $\sigma_{z}$ is given by $p(m|\sigma_{z})=\langle m|\rho|m\rangle$. Consider a unitary evolution of this state under $U=(\mathbb{1}-i\sigma_{Y})/\sqrt{2}$, the subsequent measurement of spin in the $\sigma_{3}$ direction, the probability of getting the result $m$, given by $p(m)=\langle m|U^{\dagger}\rho U|m\rangle$ is the probability of getting the result $m$ after a $\sigma_{1}$ measurement on $\rho$. This is because $U|m\rangle$ is the eigenvector of $\sigma_{x}$ with eigenvalue $m$. In general, if one wants to measure the observable $\vec{\sigma}.\vec{n}$ where $\vec{n}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)$, it is equivalent to a $\sigma_{z}$ measurement of the state $U\rho U^{\dagger}$ where $U=e^{-i(\theta/2)\sigma.\hat{n}_{\perp}}$ is unitary with $\hat{n}_{\perp}=(-\sin\phi,\cos\phi,0)$. This is possible due to the following identity: $e^{-i(\theta/2)\sigma.\hat{n}_{\perp}}\sigma_{z}e^{i(\theta/2)\sigma.\hat{n}_{\perp}}=\vec{\sigma}.\vec{n}$  . In the Stern-Gerlach setting, the interaction Hamiltonian is given by $H=-\vec{\mu}.\vec{B}$, where $\vec{\mu}=\gamma\vec{\sigma}$ is the magnetic moment, $\gamma$ is the gyromagnetic factor, and $\vec{B}$ is the magnetic field. The evolution $U$ is carried out by fixing the this evolution is carried out by the magnetic field $\vec{B}=\theta\hat{n}_{\perp}/(\gamma t)$.
Given the data $\{m,p(m|\hat{n}_{i})\}$, one can construct the density matrix using the Born rule. The probabilities are related to the state in the following way: $p(m|n_{3})=\langle m|\rho|m\rangle$, such that $\sum mp(m|n_{3})=tr(\rho\sigma_{3})$ is the expectation value of the $\sigma_{3}$ measurement. Fixing the magnetic field such that $\gamma Bt\vec{\sigma}.\hat{n}=\pi/4\sigma_{1}$  and then passing the particle through a secondary magnetic field in $z-$direction, one measures $\sigma_{2}$. The probabilities are given by $p(m|n_{1})=\langle m|U_{1}\rho U_{1}^{\dagger}|m\rangle$ where $U_{1}=e^{-i\pi/4\sigma_{2}}$. Similarly, for $U_{2}=e^{i\pi/4\sigma_{1}}$, we have $p(m|n_{2})=\langle m|U_{2}\rho U_{2}^{\dagger}|m\rangle$.